# Quantum geometry and low-frequency optical conductivity of topological nodal planes

### v2

@@ Abstract
Nodal planes, two-dimensional symmetry-enforced band crossings, can carry a topological charge, similar to Weyl points. While the transport properties of Weyl points are well understood, those of nodal planes remain largely unexplored. These properties are influenced not only by the Berry curvature, but also by other quantum geometric quantities. In this work we study the quantum geometry — specifically the Berry curvature and quantum metric — and the linear optical conductivity of topological nodal planes. We introduce a low-energy model and investigate its low-frequency optical responses to determine the unique signatures of topological nodal planes. By comparing these findings to the optical response in a tight-binding model with a topological nodal plane, we observe consistent low-frequency behavior with a cubic power law. This paves the way for the experimental detection of topological nodal planes through optical conductivity measurements.
@@ I Introduction
Optical response measurements are particularly useful because they can be performed over a wide range of frequencies, temperatures, and external conditions, and do not require big samples or large-scale measurement facilities. The versatility of optical measurements allows for the exploration of various physical phenomena and a fast screening of a large number or samples, potentially enabling the discovery of new topological systems. Therefore, a better understanding of the optical responses does not only deepen our fundamental knowledge of topological semimetals, but is also useful to screen material candidates and for the development of novel applications based on topological responses.
While the optical responses of Weyl and Dirac points are relatively well understood, those of nodal planes remain mostly unknown. This is in part because topological semimetals with nodal planes, i.e., degeneracies on two-dimensional (2D) planes in the BZ, have only recently gained attention [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 11, 12, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. The advantage compared to topological nodal features of lower dimension, like Weyl points, is that nodal planes cover a larger part of the Brillouin zone and, if dispersive, a large energy range in the spectrum. This means that less fine tuning is needed in order for them to be part of the Fermi surface. Recently, a large Nernst response of (weakly gapped) topological nodal planes has been detected [37]. This raises the question whether nodal plane materials generically show particularly large or characteristic optical responses, which can be traced back to their nontrivial topology and characteristic quantum geometry.
In this work, we provide a simple, but generic, low-energy continuum model of a topological nodal plane which can be used to tackle this question. We first discuss the properties of the band structure of the model before analyzing its quantum geometry, i.e., the Berry curvature and quantum metric. The linear optical conductivity is directly determined by the quantum metric. We analytically calculate its component perpendicular to the nodal plane and find a generic $\omega^{3}$ power law for low frequencies $\omega$. To verify the generality of this low-energy power law, we then numerically compute the optical conductivity of a generic tight-binding model with a nodal plane and find very good agreement with the analytics.
@@ II Low-energy model of a topological nodal plane
A stable nodal degeneracy on a 2D surface, called nodal plane, in the Brillouin zone (BZ) can be enforced by the combined symmetry $\tilde{\mathcal{C}}_{2}^{a}\mathcal{T}$. It consists of a two-fold screw rotation $\tilde{\mathcal{C}}_{2}^{a}$ around the axis $a$ with a non-symmorphic lattice translation by $\frac{1}{2}$ of the unit cell along the rotation axis and time-reversal symmetry $\mathcal{T}$ [23, 42]. One finds that $\tilde{\mathcal{C}}_{2}^{a}\mathcal{T}$ squares to $-1$ at $k_{a}=\pi$ leading to a two-fold degeneracy for all points on the $k_{a}=\pi$ plane due to Kramers theorem.
Figure 1: (a), (d), (g) Band structure of the low-energy model, Eq. (1), at $k_{y}=0$ for $a=1$, $a=0$, and $a=-1$, respectively. The position of the nodal plane is marked by a gray plane indicating its extension in the $k_{y}$-direction. The distribution of Berry curvature $\Omega^{xy}$ in the model for the three cases is shown on (b), (e), (h) and the distribution of quantum metric $g^{zz}$ in (c), (f), (i). Note that the distribution of quantum geometry only depends on $k_{z}$ and the radial distance $k_{x}^{2}+k_{y}^{2}$ which is used in these figures. For the figures showing the distribution of quantum geometry, the scales were cut in order for the details to be visible. The fact that these quantities are diverging in the cases of $a=0$ (and $a=-1$) is indicated in the insets of (e) and (f) (and hold also for (h) and (i)).
For the construction of a low-energy model of a topological nodal plane, the only restriction is that it must be symmetric under this $\tilde{\mathcal{C}}_{2}^{a}\mathcal{T}$ symmetry, where we choose $a=z$ for the rotation axis, leading to a nodal plane at $k_{z}=0$. Up to cubic order in $k$, we obtain the Hamiltonian
@@
which fulfills these requirements and exhibits a flat topological nodal plane at $k_{z}=0$ with charge $\nu=\text{sign}(a)$. The Pauli matrices are denoted by $\sigma_{i}$ and the parameter $\varepsilon$ determines the dispersion of the bands away from the nodal plane, i.e., in the direction perpendicular to the nodal plane. In particular, close to the nodal plane and for $a\neq 0$ the dispersion is linear with the slope $\varepsilon a$. Since the charge of the nodal plane depends on $a$, by varying it, one can observe a phase transition between $\nu=-1$ and $\nu=+1$, which is accompanied by the creation of two Weyl points. The case $a=0$ exhibits a special behaviour, which is discussed in more detail in the following sections. In terms of the band structure, however, it is qualitatively the same as for $a>0$. The band structure for the three cases, $a=+1$, $a=0$, and $a=-1$, are shown in Figs. 1 (a),(d),(g), respectively. Note that this low-energy model is not unique, since not all allowed cubic terms are included. It is, however, the simplest model which exhibits the generic features discussed in the following sections. An argument for this and the discussion of low-energy models of different order in $k$ are given in Appendix A.
@@ III Quantum geometry of the low-energy model
with $\mathbf{d}(\mathbf{k})=(d_{x},d_{y},d_{z})^{T},d_{i}\in\mathbb{R}$ ($i\in\{0,x,y,z\}$) and the vector of Pauli matrices $\bm{\sigma}$. The energies of the two bands are $E_{\pm}=d_{0}\pm|\mathbf{d}(\mathbf{k})|$. One can define a normalized vector $\mathbf{n}(\mathbf{k})=\frac{\mathbf{d}(\mathbf{k})}{|\mathbf{d}(\mathbf{k})|}$ with which the QGT of the two bands denoted by $\pm$ can be expressed as
$\mathcal{Q}^{ab}_{\pm}$ $=\frac{1}{4}\partial_{a}\mathbf{n}\cdot\partial_{b}\mathbf{n}\mp\frac{i}{4}\mathbf{n}\cdot(\partial_{a}\mathbf{n}\times\partial_{b}\mathbf{n})$ (3)
$=g^{ab}_{\pm}+\frac{i}{2}\Omega^{ab}_{\pm}\;,$
@@
With this, we can discuss the quantum geometry of the nodal plane low-energy model defined in Eq. (1) for the three cases of $a$ discussed above. However, for a better intuition, we will briefly review the case of a Weyl point described by the low-energy model $\mathcal{H}_{\mathrm{WP}}=\mathbf{k}\cdot\bm{\sigma}$. For this 0D nodal feature it is well known that the Berry curvature given by $\mathbf{\Omega}_{\pm}=\pm\frac{\mathbf{k}}{2(k_{x}^{2}+k_{y}^{2}+k_{z}^{2})^{\frac{3}{2}}}$ diverges at the crossing and the charge of the Weyl point can be calculated by determining the flux of Berry curvature through a sphere enclosing it. In fact, also the quantum metric diverges at that point.
@@
In Figs. 1 (b),(e),(h) we show the distribution of the Berry curvature of the entire system for the three choices of $a$ discussed in the previous section. One finds that for $a=1$ the Berry curvature has the expected jump at the nodal plane and is spread broadly in the $k_{x}^{2}+k_{y}^{2}$-plane. The charge is $\nu=+1$. The case $a=0$ is a special point since here the Berry curvature diverges at $\mathbf{k}=0$, as indicated in the inset of Fig. 1(e), which resembles the monopole behaviour of a Weyl point and, as for $a=1$, it integrates to $\nu=+1$. For $a=-1$ two additional Weyl points display the characteristic divergence of Berry curvature and the sign jump at the nodal plane is reversed, leading to a charge of $\nu=-1$. We find that the $\Omega^{xy}$ component is sufficient to study the topological properties of the nodal plane, since its behaviour is determined by the symmetry enforcing the nodal plane.
@@
which is positive definite and symmetric in $k_{z}$. The distribution of $g^{zz}$ is, as for the Berry curvature, shown for the three cases of $a$ in Figs. 1 (c),(f),(i). We find again that its distribution depends on $a$. In particular, for $a=1$, the metric is broadly distributed around the nodal plane. For $a=0$ the metric collapses to $\mathbf{k}=0$, as indicated in the inset of Fig. 1(f), showing the discussed Weyl point behaviour. For $a=-1$ the two emerging Weyl points again dominate the metric distribution. This shows that the distribution roughly follows the one of the Berry curvature. Therefore, topological nodal features, such as Weyl points and nodal planes can be seen as the source of non-trivial quantum geometry. Importantly, this non-trivial behaviour is not restricted to signatures of the Berry curvature or quantum metric. Rather, we expect that, depending on the symmetries of the system, this will also show up in higher-order quantum geometric quantities, such as the quantum connection [48, 67, 49, 52, 68]. The discussion of this is, however, beyond the scope of this work.
@@ IV Linear optical conductivity
To evaluate this for the nodal plane low-energy model at $T=0$, we consider the cases $a\geq 0$ and $a<0$ separately. In Fig. 2, the results for the different values of parameter $a$ are shown. The full analytical expression for $a\geq 0$ is given in Appendix C and, in the following, the leading-order behaviour is discussed. We find that for $a>0$ the leading order in $\omega$ in the Taylor expansion of the conductivity is cubic, i.e.,
@@
Figure 2: $zz$-component of the interband linear conductivity, $\sigma^{zz}$, for the low-energy model, Eq. (1), as a function of frequency $\omega$ for three different values of $a$ with $\varepsilon=1$.
@@
Figure 3: (a) Band structure of the modified tight-binding model introduced in [42] with a topological nodal plane at $k_{z}=\pi$ and a two Weyl points at the K points. The opacity of the bands indicates the areas of large (diverging) $zz$-component of the quantum metric. The inset shows the hexagonal BZ with the high-symmetry points. (b) $zz$-component of the linear optical conductivity for the tight-binding (TB) model calculated numerically using a phenomenological broadening of $\Gamma=0.02$. A fit of the form $x^{\alpha}$ in the low-frequency region ($\omega\leq 0.5$) yields the power-law behaviour $\alpha=3.07\pm 0.01$ and is indicated with a pink line. The analytical result for the low-energy model with the parameters of the tight-binding model is shown for low frequencies with the dashed line (LE). The inset shows the low-frequency behaviour using logarithmic scales on both axes.
@@
While the results for $a\leq 0$ are interesting and in agreement with an intuitive understanding of these topological features, the $\omega^{3}$ dependence for $a>0$, i.e., the case with a single topological nodal plane, is completely novel and, due to the enforcing symmetry that impacts the $zz$-component of the quantum metric, expected to be a characteristic signature of a topological nodal plane. Higher-order terms in $k$ in the low-energy description can only impact higher-order terms in the $\omega$-expansion. This means that, if a material has a single topological nodal plane at the Fermi energy (without additional topological features nearby), this characteristic frequency dependence can be used to detect the nodal plane, similarly to the detection of Weyl points through their linear behaviour [15]. In order to support this statement and as a first step to investigate whether this characteristic behaviour can be observed in more complicated systems, we study the optical response of a generic tight-binding model with a topological nodal plane in the following section. Note that this characteristic signature is not restricted to topologically charged nodal planes. In Appendix E, we introduce a minimal low-energy model of a nodal plane with vanishing Berry curvature but finite quantum metric and show that also in this case the leading-order contribution in the linear optical conductivity component perpendicular to the nodal plane is $\omega^{3}$.
V Optical conductivity of a tight-binding model

### v3

# Quantum geometry and low-frequency optical conductivity of nodal planes

@@ Abstract
Nodal planes, two-dimensional symmetry-enforced band crossings, can carry a topological charge, similar to Weyl points. While the transport properties of Weyl points are well understood, those of nodal planes remain largely unexplored. These properties are influenced not only by the Berry curvature, but also by other quantum geometric quantities. In this work we study the quantum geometry — specifically the Berry curvature and quantum metric — and the linear optical conductivity of topological nodal planes. We introduce a low-energy model and investigate its low-frequency optical responses to determine the unique signatures of nodal planes. By comparing these findings to the optical response in a tight-binding model with a topological nodal plane, we observe consistent low-frequency behavior with a cubic power law. This paves the way for the experimental detection of nodal planes through optical conductivity measurements for which we suggest suitable materials, most promisingly the material group XXMo3S3 ($X$ = Rb, K, Cs).
@@ I Introduction
Optical response measurements are particularly useful because they can be performed over a wide range of frequencies, temperatures, and external conditions, and do not require big samples or large-scale measurement facilities. The versatility of optical measurements allows for the exploration of various physical phenomena and a fast screening of a large number of samples, potentially enabling the discovery of new topological systems. Therefore, a better understanding of the optical responses does not only deepen our fundamental knowledge of topological semimetals, but is also useful to screen material candidates and for the development of novel applications based on topological responses.
While the optical responses of Weyl and Dirac points are relatively well understood, those of nodal planes remain mostly unknown. This is in part because topological semimetals with nodal planes, i.e., degeneracies on two-dimensional (2D) planes in the Brillouin zone (BZ), have only recently gained attention [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 11, 12, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. The advantage compared to topological nodal features of lower dimension, like Weyl points, is that nodal planes cover a larger part of the BZ and, if dispersive, a large energy range in the spectrum. This means that less fine tuning is needed in order for them to be part of the Fermi surface. Recently, a large Nernst response of (weakly gapped) topological nodal planes has been detected [37]. This raises the question whether nodal plane materials generically show particularly large or characteristic optical responses, which can be traced back to their nontrivial topology and characteristic quantum geometry.
In this work, we provide a simple, but generic, low-energy continuum model of a topological nodal plane which can be used to tackle this question. We first discuss the properties of the band structure of the model before analyzing its quantum geometry, i.e., the Berry curvature and quantum metric. The linear optical conductivity is directly determined by the quantum metric. We analytically calculate its component perpendicular to the nodal plane and find a generic $\omega^{3}$ power law for low frequencies $\omega$. To verify the generality of this low-energy power law, we then numerically compute the optical conductivity of a generic tight-binding model with a nodal plane and find very good agreement with the analytics. Finally, we discuss requirements for the experimental detection of the signature of nodal planes in optical conductivity measurements and list suitable materials.
@@ II Low-energy model of a topological nodal plane
A stable nodal degeneracy on a 2D surface, called nodal plane, in the BZ can be enforced by the combined symmetry $\tilde{\mathcal{C}}_{2}^{a}\mathcal{T}$. It consists of a two-fold screw rotation $\tilde{\mathcal{C}}_{2}^{a}$ around the axis $a$ with a non-symmorphic lattice translation by $\frac{1}{2}$ of the unit cell along the rotation axis and time-reversal symmetry $\mathcal{T}$ [23, 42]. One finds that $\tilde{\mathcal{C}}_{2}^{a}\mathcal{T}$ squares to 1-1 at $k_{a}=\pi$ leading to a two-fold degeneracy for all points on the $k_{a}=\pi$ plane due to Kramers theorem.
Figure 1: (a), (d), (g) Band structure of the low-energy model, Eq. (1), at $k_{y}=0$ with $\varepsilon=1$ for $a=1$, $a=0$, and a=1a=-1, respectively. The position of the nodal plane is marked by a gray plane indicating its extension in the $k_{y}$-direction. The distribution of Berry curvature $\Omega^{xy}$ in the model for the three cases is shown in (b), (e), (h) and the distribution of quantum metric $g^{zz}$ in (c), (f), (i). Note that the distribution of quantum geometry only depends on $k_{z}$ and the radial distance $k_{x}^{2}+k_{y}^{2}$ which is used in these figures. For the figures showing the distribution of quantum geometry, the scales were cut in order for the details to be visible. The fact that these quantities are diverging in the cases of $a=0$ (and a=1a=-1) is indicated in the insets of (e) and (f) (and hold also for (h) and (i)).
For the construction of a low-energy model of a topological nodal plane, the only restriction is that it must be symmetric under this $\tilde{\mathcal{C}}_{2}^{a}\mathcal{T}$ symmetry, where we choose $a=z$ for the rotation axis, leading to a nodal plane at $k_{z}=0$. The $\tilde{\mathcal{C}}_{2}^{z}\mathcal{T}$ symmetry acts on the Pauli matrices as $\sigma_{i}\rightarrow-\sigma_{i}$ for $i\in\{x,y,z\}$ and on the momenta as $k_{x/y}\rightarrow k_{x/y}$ and $k_{z}\rightarrow-k_{z}$. Therefore, for the symmetry to be fulfilled, every term in the Hamiltonian, except for trivial terms $\sim\mathbb{1}$, has to be odd in $k_{z}$. Up to cubic order in $k$, we obtain the Hamiltonian
@@
which fulfills these requirements and exhibits a flat topological nodal plane at $k_{z}=0$ with charge $\nu=\text{sign}(\varepsilon a)$. The Pauli matrices are denoted by $\sigma_{i}$ and the parameter $\varepsilon$ determines the dispersion of the bands away from the nodal plane, i.e., in the direction perpendicular to the nodal plane. In particular, close to the nodal plane and for $a\neq 0$ the dispersion is linear with the slope $\varepsilon a$. We set $\varepsilon=1$ for this discussion. Since the charge of the nodal plane depends on $a$, by varying it, one can observe a phase transition between $\nu=+1$ and $\nu=-1$, which is accompanied by the creation of two Weyl points. The case $a=0$ exhibits a special behaviour, which is discussed in more detail in the following sections. In terms of the band structure, however, it is qualitatively the same as for $a>0$. The band structure for the three cases, $a=+1$, $a=0$, and a=1a=-1, are shown in Figs. 1 (a),(d),(g), respectively. Note that this low-energy model is not unique, since not all allowed cubic terms are included. It is, however, the simplest model which exhibits the generic features discussed in the following sections. An argument for this and the discussion of low-energy models of different order in $k$ are given in Appendix A.
@@ III Quantum geometry of the low-energy model
with $\mathbf{d}(\mathbf{k})=(d_{x},d_{y},d_{z})^{T},d_{i}\in\mathbb{R}$ ($i\in\{0,x,y,z\}$) and the vector of Pauli matrices $\bm{\sigma}$. The energies of the two bands are $E_{\pm}=d_{0}(\mathbf{k})\pm|\mathbf{d}(\mathbf{k})|$. One can define a normalized vector $\mathbf{n}(\mathbf{k})=\frac{\mathbf{d}(\mathbf{k})}{|\mathbf{d}(\mathbf{k})|}$ with which the QGT of the two bands denoted by $\pm$ can be expressed as
$\mathcal{Q}^{ab}_{\pm}$ $=\frac{1}{4}\partial_{a}\mathbf{n}\cdot\partial_{b}\mathbf{n}\pm\frac{i}{4}\mathbf{n}\cdot(\partial_{a}\mathbf{n}\times\partial_{b}\mathbf{n})$ (3)
$=g^{ab}_{\pm}-\frac{i}{2}\Omega^{ab}_{\pm}\;,$
@@
With this, we can discuss the quantum geometry of the nodal plane low-energy model defined in Eq. (1) for the three cases of $a$ discussed above. However, for a better intuition, we will briefly review the case of a Weyl point described by the low-energy model $\mathcal{H}_{\mathrm{WP}}=\mathbf{k}\cdot\bm{\sigma}$. For this 0D nodal feature it is well known that the Berry curvature given by $\mathbf{\Omega}_{\pm}=\mp\frac{\mathbf{k}}{2(k_{x}^{2}+k_{y}^{2}+k_{z}^{2})^{\frac{3}{2}}}$ diverges at the crossing and the charge of the Weyl point can be calculated by determining the flux of Berry curvature through a sphere enclosing it. In fact, also the quantum metric diverges at that point.
@@
In Figs. 1 (b),(e),(h) we show the distribution of the Berry curvature of the entire system for the three choices of $a$ discussed in the previous section. One finds that for $a=1$ the Berry curvature has the expected sign jump at the nodal plane and is spread broadly in the $k_{x}^{2}+k_{y}^{2}$-plane. The charge is $\nu=+1$. The case $a=0$ is a special point since here the Berry curvature diverges at $\mathbf{k}=0$, as indicated in the inset of Fig. 1(e), which resembles the monopole behaviour of a Weyl point and, as for $a=1$, it integrates to $\nu=+1$. For a=1a=-1 two additional Weyl points display the characteristic divergence of Berry curvature and the sign jump at the nodal plane is reversed, leading to a charge of $\nu=-1$. We find that the $\Omega^{xy}$ component is sufficient to study the topological properties of the nodal plane, since its behaviour is determined by the symmetry enforcing the nodal plane.
@@
which is positive definite and symmetric in $k_{z}$. The distribution of $g^{zz}$ is, as for the Berry curvature, shown for the three cases of $a$ in Figs. 1 (c),(f),(i). We find again that its distribution depends on $a$. In particular, for $a=1$, the metric is broadly distributed around the nodal plane. For $a=0$ the metric collapses to $\mathbf{k}=0$, as indicated in the inset of Fig. 1(f), showing the discussed Weyl point behaviour. For a=1a=-1 the two emerging Weyl points again dominate the metric distribution. This shows that the distribution roughly follows the one of the Berry curvature. Therefore, topological nodal features, such as Weyl points and nodal planes, can be seen as the source of non-trivial quantum geometry. Importantly, this non-trivial behaviour is not restricted to signatures of the Berry curvature or quantum metric. Rather, we expect that, depending on the symmetries of the system, this will also show up in higher-order quantum geometric quantities, such as the quantum connection [48, 67, 49, 52, 68]. The discussion of this is, however, beyond the scope of this work.
@@ IV Linear optical conductivity
As discussed for the quantum geometry, we will focus on the optical conductivity component perpendicular to the nodal plane $\sigma^{zz}$, which is determined by the quantum metric component $g^{zz}$. Only this component is restricted by the $\tilde{\mathcal{C}}_{2}^{z}\mathcal{T}$ symmetry, and therefore expected to show generic features. To evaluate this for the nodal plane low-energy model at $T=0$, we consider the cases $a\geq 0$ and $a<0$ separately. In Fig. 2, the results for the different values of parameter $a$ are shown. The full analytical expression for $a\geq 0$ is given in Appendix C and, in the following, the leading-order behaviour is discussed. We find that for $a>0$ the leading order in $\omega$ in the Taylor expansion of the conductivity is cubic, i.e.,
@@
Figure 2: zzzz-component of the interband linear conductivity, $\sigma^{zz}$, for the low-energy model, Eq. (1), as a function of frequency $\omega$ for three different values of $a$ with $\varepsilon=1$.
@@
Figure 3: (a) Band structure of the modified tight-binding model introduced in [42] with a topological nodal plane at $k_{z}=\pi$ and a two Weyl points at the K points. The opacity of the bands indicates the areas of large (diverging) zzzz-component of the quantum metric. The inset shows the hexagonal BZ with the high-symmetry points. (b) zzzz-component of the linear optical conductivity for the tight-binding (TB) model calculated numerically using a phenomenological broadening of $\Gamma=0.02$. A fit of the form $x^{\alpha}$ in the low-frequency region ($\omega\leq 0.5$) yields the power-law behaviour $\alpha=3.07\pm 0.01$ and is indicated with a pink line. The analytical result for the low-energy model with the parameters of the tight-binding model is shown for low frequencies with the dashed line (LE). The inset shows the low-frequency behaviour using logarithmic scales on both axes.
@@
While the results for $a\leq 0$ are interesting and in agreement with an intuitive understanding of these topological features, the $\omega^{3}$ dependence for $a>0$, i.e., the case with a single topological nodal plane, is completely novel and, due to the enforcing symmetry that impacts the $zz$-component of the quantum metric, expected to be a characteristic signature of a topological nodal plane. Higher-order terms in $k$ in the low-energy description can only impact higher-order terms in the $\omega$-expansion. This means that, if a material has a single topological nodal plane at the Fermi energy (without additional topological features nearby), this characteristic frequency dependence can be used to detect the nodal plane, similarly to the detection of Weyl points through their linear behaviour [15].
V Trivial nodal planes and arbitrary dispersions
The characteristic cubic frequency dependence is not restricted to topologically charged nodal planes. A minimal low-energy model of a trivial nodal plane is given by
$g^{zz}_{\pm}=\frac{2k_{z}^{2}}{(3+2k_{z}^{2}+k_{z}^{4})^{2}}\;.$ (12)
As for the topological nodal plane model, we calculate the linear optical conductivity, where in this case only the component perpendicular to the nodal plane $\sigma^{zz}$ is non-zero. The full expression is given in Appendix E and we obtain again a cubic power law in the low-frequency expansion given by $\sigma_{zz}=\frac{\pi}{(2\pi)^{3}}\frac{\omega^{3}}{54\sqrt{3}}+\mathcal{O}\left(\omega^{5}\right)$.
Furthermore, we show in Appendix B that our findings also hold when including arbitrary (symmetry-allowed) dispersion terms in the Hamiltonian, i.e., $\mathcal{H}_{\text{disp}}=\mathcal{H}+d_{0}(\mathbf{k})\mathbb{1}$, since they enter the optical conductivity only through the Fermi-Dirac distribution. With this, also a variation of the chemical potential, corresponding to a constant $d_{0}$, cannot change this behaviour. This discussion holds on the level of low-energy models, while in tight-binding models other band features and the band width have to be taken into account. In particular, if other topological nodal features such as Weyl points occur in the vicinity of the Fermi energy, these can dominate the low-frequency optical response.
VI Optical conductivity of a tight-binding model
In order to study, whether the characteristic low-frequency behaviour is visible also in more complicated systems, we consider a generic tight-binding model exhibiting a single nodal plane. Specifically, we examine a two-band tight-binding model for a hexagonal space group, which was introduced in [42], see Appendix D. We modify its dispersion by adding the term $\cos(k_{z})\mathbb{1}$ to increase the energy separation of the nodal plane and the Weyl points. This is crucial in order to study the signatures of the nodal plane alone, since, as shown in Sec. IV, Weyl points close to the Fermi level will likely dominate the optical response. We choose the chemical potential such that a broad region of the nodal plane lies close to the Fermi energy. The resulting band structure is shown in Fig. 3(a). It has a nodal plane at $k_{z}=\pi$, which carries a topological charge $\nu=2$, and two Weyl points at the K points, which compensate the nodal-plane charge.